The hypotenuse of a right triangle is 25 cm long. one leg is 17 cm longer than the other. Find the length of the shorter leg.
___cm
Thank you

Answer by fractalier(2101)

Let the shorter leg be x. The longer is then x + 17. Using the Pythagorena Theorem, we have
x^2 + (x + 17)^2 = 25^2
x^2 + x^2 + 34x + 289 = 625
2x^2 + 34x - 336 = 0
x^2 + 17x - 168 = 0
(x + 24)(x - 7) = 0
x = -24 or x = 7
but distance cannot be negative so
x = 7